OpenStudy (anonymous):

Landing with a speed of 75.6, and traveling due south, a jet comes to rest in 933. Assuming the jet slows with constant acceleration, find the magnitude and direction of its acceleration.

OpenStudy (anonymous):

Ok, I got this problem wrong but I'm not sure why. I'll post my work.

OpenStudy (anonymous):

*75.6m/s *933m

OpenStudy (anonymous):

t=933/75.6 => t=12.3 s a=v/t => 75.6/12.3= 6.14m/s^2

OpenStudy (anonymous):

The accelartion is negative

OpenStudy (anonymous):

Yeah, I actually had -6.14 m/s^2 and it was wrong.

OpenStudy (anonymous):

\[0^{2}=75.6^{2}+2a933\]

OpenStudy (anonymous):

Can you explain that formula?

OpenStudy (anonymous):

\[v ^{2}=v _{0}^{2}+2ad\]